OpenStudy (anonymous):

find two consecutive integers that have a sum of -67

jimthompson5910 (jim_thompson5910):

if x is the first integer, then x+1 is the next consecutive integer

jimthompson5910 (jim_thompson5910):

They must add to -67, so (x) + (x+1) = -67 x + x + 1 = -67 2x + 1 = -67 Solve for x. Then use this to find x and x+1, which are the two consecutive integers.